OpenStudy (anonymous):

Find the inverse of the function below. Graph the function below and the inverse function. Determine the domain, range, and asymptotes of the function below and the inverse function. Please show your work. f(x) = 3 ln(4-x)+1

OpenStudy (anonymous):

y = 3 ln(4-x)+1, solving for x (y-1)/3 = ln(4-x) taking exponential power on both sides we get e^[(y-1)/3] = e^ln(4-x) e^[(y-1)/3] = (4-x) x=4-e^[(y-1)/3] interchange x to y y =4-e^[(x-1)/3] answer....inverse function from the function f(x) = 3 ln(4-x)+1 vertical retriceymptote x=4 range(-∞,∞) Can anyone show me how to graph this??!